# An Analysis of Concept Bottleneck Models: Measuring, Understanding, and Mitigating the Impact of Noisy Annotations

### v1

@@ 1 Introduction
Despite a decade of breakthroughs in deep learning, the inner-workings of end-to-end neural networks remain largely elusive [34, 23, 33]. This lack of transparency stems from their reliance on complex, high-dimensional internal representations, which offer limited insight into their decision-making process [36, 7]. Concept bottleneck models (CBMs) are designed to address this opacity by first predicting a set of human-interpretable concepts, which serve as an intermediate representation for deriving the final decision [18, 2, 37]. For instance, when it comes to the bird species classification problem [52], it first predicts the concepts of the bird such as tail shape or body color, and then, infers the bird class based on these attributes, making the decision more understandable (e.g., “this bird is Heermann Gull because it is gray and black in color with an orange beak!”).
Formally, a CBM is trained on triplets of input $x\in\mathbb{R}^{d}$, binary concept vectors $c\in\{0,1\}^{k}$, and output $y$. The concept predictor $g:x\to c$ maps each input to its $k$-dimensional concept vector, and the output predictor $f:c\to y$ consumes those concepts to produce $y$ (see Figure 1(a)). This two-stage process offers a practical advantage, which is, domain experts can step in and edit the predicted concepts at inference, a process known as concept-level intervention [18]. For example, changing ‘has stripe’ from yes to no, while keeping other concepts such as ‘number of legs’, can shift the prediction from zebra to horse without retraining and can recover much of the lost accuracy. Interventions further support counterfactual analysis [51], allowing users to ask “what if” scenarios by manipulating concepts rather than raw pixels. Such human-in-the-loop interactivity is central to the appeal of CBMs and underscores the need to safeguard concept fidelity.
@@
Specifically, our study reveals that noise significantly compromises key characteristics of CBMs, i.e., interpretability, intervention effectiveness, as well as prediction performance (see Figure 1 for basic results). We discover that this degradation is primarily driven by a specific set of concepts that are highly susceptible to noise, whose accuracy drop largely accounts for overall performance declines. To mitigate these negative effects, we propose two strategies. First, we utilize sharpness-aware minimization (SAM) [9] during training to build robustness, particularly for noise-sensitive concepts. Second, observing that susceptible concepts often have high predictive uncertainty, quantified by entropy [40], we adopt an uncertainty-based intervention strategy at inference [20, 19, 42]. This strategy prioritizes correcting high-uncertainty concepts, enabling targeted interventions that significantly improve overall task performance. Extensive empirical results validate the effectiveness of this combined approach, consistently recovering the reliability of CBMs despite the presence of noise.
@@ 2 Measuring Impact
In this section, we systematically measure the impact of noise on CBMs across three key dimensions: prediction performance, interpretability, and intervention effectiveness. We begin by outlining our experimental setup (Section 2.1), followed by an evaluation of the noise effects on prediction performance (Section 2.2), interpretability (Section 2.3), and intervention effectiveness (Section 2.4).
@@ 2.3 Interpretability
Next, we assess the effect of noise on interpretability. As shown in Figure 2(c), concept prediction accuracy declines from 96.5% in clean conditions to 85.4% with 40% noise on the CUB dataset. Despite appearing modest, the 11.1% drop notably compresses the margin over random guessing in binary classification tasks, where minor concept prediction errors can compound and severely impair target prediction in CBMs. Beyond accuracy, interpretability critically relies on semantic coherence of learned representations. Thus, we measure interpretability using the concept alignment score (CAS) [59], quantifying semantic alignment between learned representations and ground-truth concepts (see Figure 2(d)). We find CAS decreases substantially with increasing noise, e.g., for each additional 10% increment of noise, CAS drops by approximately 7% on the CUB dataset, indicating severe degradation in interpretability.
Qualitative t-SNE [49] visualizations further illustrate this interpretability decline (see Figure 4). Clean embeddings for the specific concept ‘blue upperparts’ show distinct, semantically meaningful clusters corresponding to active (i.e., red circles) and inactive (i.e., blue circles) concept instances. However, increasing noise levels result in ambiguous, less separable clusters. Thus, noise compromises the explanatory power and reliability of concepts, diminishing interpretability. Additional examples are provided in Section D.11.
@@ 2.4 Intervention Effectiveness
Finally, we investigate how noise limits the utility of post-hoc interventions, utilizing three concept selection strategies proposed by Shin et al. [42]: (i) Random, which samples concepts uniformly [18]; (ii) UCP (uncertainty of concept prediction), which ranks concepts by predictive uncertainty; (iii) CCTP (contribution of concept on target prediction), which selects concepts exerting the greatest influence on the target prediction. Implementation specifics are detailed in Section C.5.
As shown in Figure 5(a)5(c), the benefits of intervention are increasingly constrained as noise levels rise (i.e., the accuracy curves translate downward along the y-axis). Under the Random strategy, correcting every concept in a model trained with clean supervision yields near-optimal performance at 99.8% accuracy. However, this upper bound progressively declines to 98.4%, 95.7%, 89.4%, and 53.1% as the noise level increases from 10% to 40%. Notably, even exhaustive intervention at 40% noise fails to recover the performance of the clean model without any intervention. Similar trends are observed for the advanced intervention methods (e.g., UCP and CCTP), indicating that the errors induced by noisy supervision in CBMs cannot be rectified through post-hoc correction alone.
To further evaluate robustness, we deliberately assign incorrect values to selected concepts, emulating the inevitable inaccuracies of expert intervention. As shown in Figure 5(d)5(e), CBMs trained on clean data tolerate a small number of such mistakes with only marginal performance decline. Under noise, however, even slight intervention errors precipitate pronounced additional loss, revealing a fragility that limits the practical utility of concept-level correction in realistic human-expert workflows.
@@ 3 Understanding Mechanisms
In this section, we further analyze how noise affects the internal mechanisms of CBMs, separately examining the concept predictor $g$ and the target predictor $f$. We begin with the observation that noise induces non-uniform degradation of individual concept accuracies (Section 3.1). We then assess how noise disrupts the alignment between concepts representations and target predictions (Section 3.2).
@@ 3.1 Non-Uniform Degradation of Concept Accuracy
As shown in Figure 6, although noise is introduced uniformly across all concepts, accuracy degradation is notably uneven. While the majority of concepts remain comparatively resilient, a distinct minority suffers a disproportionately large decline. We designate this minority as the susceptible concept set, defined as
@@
To better understand this phenomenon, we plot the concept frequency distributions using kernel density estimation (KDE) [4] in Figure 7. In the clean dataset, concept occurrences are imbalanced, with many low-frequency and a few high-frequency concepts. As noise increases, however, this distribution becomes more uniform; at 40% noise, most concepts occur roughly 2,000 times. This shift may reduce the ability of model to learn rare but informative concepts, as the effective signal-to-noise ratio for such concepts deteriorates. Identifying and understanding these vulnerable concepts is essential, as their degradation can disproportionately affect downstream performance, which we explore in the next section.
@@ 3.2 Disruption in Representation Alignment
We next examine how noise during training reshapes the correspondence between learned representations and their associated target classes. In a noise-free setting we first identify the five most influential representation dimensions (i.e., concepts) for the class Le Conte Sparrow. These dimensions align with meaningful visual attributes such as ‘white upperparts’ and ‘grey back’. Their importance is quantified by the absolute magnitude of the weights in the linear target predictor $f$, a widely used proxy for feature saliency [24, 21, 5]. We then trace how these saliency values change as the proportion of corrupted labels increases.
As illustrated in Figure 8(b), when the noise level is moderate (i.e., 20%) the relative ordering of the predictive dimensions is largely preserved. At a higher noise level (i.e., 40%), however, their saliency declines markedly. Informative dimensions lose influence, whereas irrelevant or spurious ones gain disproportionately large weights (see Figure 8(c)). Specifically, dimensions linked to ‘grey back’ or ‘yellow upper tail’ receive minimal weight, while a related attribute such as ‘orange upper tail’ is assigned a negative weight. These shifts indicate a pronounced misalignment between the learned representations and their intended semantics.
We further relate this effect to concept-level susceptibility. In Figure 6, its five most influential dimensions, indicated with red arrows, coincide with the concepts most susceptible to noise. The same pattern emerges across the full dataset. Among 200 classes, 189 display an exact overlap between the five most influential dimensions and the noise-susceptible set. This widespread overlap exposes a critical weakness, i.e., when concepts that are noise-sensitive coincide with distorted representations, performance degrades sharply. Thus, the intersection of susceptibility and disrupted alignment emerges as a principal source of instability, emphasizing the need for robust training methods to prevent representation collapse under noisy supervision.
@@ 4 Mitigating Effects
We have shown that noise severely degrades the key aspects of CBMs, primarily due to a subset of concepts (i.e., susceptible concept set) that are especially vulnerable to noisy supervision and frequently coincide with representational misalignments. To address this, we introduce two mitigation strategies that operate at different stages of the model lifecycle: during training and inference. First, we incorporate sharpness-aware minimization (SAM[9] at training time to improve the prediction accuracy of susceptible concepts (Section 4.1). Then, at inference time, we apply an uncertainty-based intervention strategy that prioritizes corrections on unreliable concept predictions (Section 4.2).
@@ 4.1 Training-Time Robustness via SAM
where the inner maximization is approximated by $\widetilde{\bm{\varepsilon}}(\bm{w})=\rho\frac{\nabla_{\bm{w}}\ell_{\mathrm{train}}(\bm{w})}{\|\nabla_{\bm{w}}\ell_{\mathrm{train}}(\bm{w})\|_{2}}$. Gradient updates are then computed at $\bm{w}+\widetilde{\bm{\varepsilon}}(\bm{w})$. By steering solutions toward flatter optima, SAM attenuates the performance degradation caused by noisy labels and other small perturbations [9, 1].
@@
Base 96.52$\pm$0.00.$0$
$\gamma=0.0$ SAM 97.19$\pm$$0.1$
Base 91.63$\pm$$0.0$ 66.45$\pm$$1.0$ 99.02$\pm$$0.0$
$\gamma=0.2$ SAM 92.54$\pm$$0.1$ 70.30$\pm$$0.7$ +3.85 98.93$\pm$$0.1$
Base 85.42$\pm$$0.0$ 39.58$\pm$$3.0$ 98.61$\pm$$0.2$
$\gamma=0.4$ SAM 86.31$\pm$$0.1$ 43.65$\pm$$1.2$ +4.07 98.64$\pm$$0.3$
@@
We compare SAM-trained CBMs with SGD-trained ones (i.e., Base) across concept accuracy, task accuracy, and CAS on the CUB and AwA2 datasets under varying noise levels. As shown in Table 1, SAM consistently improves all metrics across noise conditions. Notably, even modest gains in concept accuracy lead to substantial improvements in downstream task performance. For instance, under 40% noise in AwA2, a 0.60% increase in concept accuracy leads to a 4.68% gain in task accuracy. Previous work [1] attributed the robustness of SAM against noise to its $\ell_{2}$-regularization on final layer weights and intermediate activations. In Appendix A, we extend the theoretical analysis of Baek et al. [1] to CBMs, demonstrating that the same regularization effect holds.
We also emphasize that SAM is particularly effective at recovering concepts in the susceptible concept set. Table 2 shows that while concept accuracy in the non-susceptible set remains stable, SAM yields substantial gains in the susceptible set across all noise levels, e.g., yielding gains of +3.85% and +4.07% under 20% and 40% noise, respectively. These findings underscore the effectiveness of SAM in selectively enhancing the learning of noise-sensitive concepts, thereby enhancing both predictive robustness and model interpretability under corrupted supervision.
@@
Figure 9: Intervention results comparing random and susceptibility-based corrections. The dotted line indicates the clean-model accuracy without intervention.
@@ 4.2 Inference-Time Robustness via Uncertainty-Guided Intervention
Although training with SAM effectively mitigates the adverse effects of noise, additional performance gains are achievable through targeted inference-time interventions. Motivated by our earlier findings that only a small subset of concepts significantly contributes to performance degradation, we prioritize corrections based on susceptibility, defined as the magnitude of accuracy decline under noise. Specifically, we select the five most susceptible concepts per class for correction. As shown in Figure 9, susceptibility-guided corrections significantly outperform random interventions across varying noise levels. Remarkably, correcting even a single highly susceptible concept can yield accuracy improvements exceeding 10% at higher noise intensities (e.g., 20% and 30%). Moreover, fewer than five targeted interventions frequently achieve or surpass the accuracy of models trained on clean data. These findings underscore that noise disproportionately affects a small subset of highly predictive and noise-sensitive concepts, and that targeted correction of these concepts substantially restores overall model performance.
@@
While susceptibility-based concept selection is a highly effective intervention strategy, it is not feasible to perform in realistic scenarios since it requires the access to a model trained with clean labels. To circumvent this challenge, we explore predictive uncertainty as a practical surrogate for susceptibility. Intuitively, concepts that contribute little to target prediction or are noise-sensitive should exhibit elevated uncertainty. To validate this hypothesis, we measure the Pearson correlation between concept-level uncertainty (quantified by the entropy of their predictive distributions [40]) and susceptibility. As depicted in Figure 10, we observe a positive correlation within the subset of susceptible concepts, whereas no consistent relationship emerges among non-susceptible concepts. These results suggest that predictive uncertainty effectively, though not definitively, reflects susceptibility induced by noise, supporting its use as a practical criterion for inference-time interventions. In Appendix B, we provide theoretical justification demonstrating that, under reasonable assumptions, uncertainty-based selection asymptotically approximates susceptibility-based concept selection.
@@
Base + Random 74.3$\pm$$0.3$ 75.6$\pm$$0.4$ 76.8$\pm$$0.5$ 50.3$\pm$$0.6$ 56.2$\pm$$0.7$ 62.1$\pm$$1.0$ 4.0$\pm$$0.6$ 4.8$\pm$$0.7$ 6.3$\pm$$0.7$
Base + Uncertainty 74.3$\pm$$0.3$ 81.8$\pm$$0.6$ 87.0$\pm$$0.7$ 50.3$\pm$$0.6$ 71.2$\pm$$1.7$ 82.0$\pm$$1.4$ 4.0$\pm$$0.6$ 6.9$\pm$$0.9$ 10.7$\pm$$1.4$
SAM + Random 79.0$\pm$$0.7$ 80.1$\pm$$0.6$ 81.1$\pm$$0.8$ 54.2$\pm$$0.6$ 59.9$\pm$$0.8$ 65.6$\pm$$0.9$ 5.0$\pm$$1.2$ 5.6$\pm$$0.8$ 7.1$\pm$$0.7$
SAM + Uncertainty 79.0$\pm$$0.7$ 86.2$\pm$$0.8$ 90.9$\pm$$0.7$ 54.2$\pm$$0.6$ 75.2$\pm$$0.7$ 85.1$\pm$$0.9$ 5.0$\pm$$1.2$ 8.0$\pm$$1.5$ 12.4$\pm$$1.7$
@@
Building on the observed correlation between predictive entropy and susceptibility, uncertainty provides a principled signal for inference-time interventions. Guided by the survey of Shin et al. [42], we employ uncertainty of concept prediction (UCP), which ranks concepts according to entropy $\mathcal{H}(\widehat{c})$ [40] and prioritize corrections of the most uncertain ones. As illustrated in Figure 11, UCP consistently outperforms baseline methods such as Random selection and CCTP, with the margin of improvement increasing at higher noise intensities. These results indicate that uncertainty-based interventions effectively approximate susceptibility-based strategies, successfully identifying and rectifying noise-sensitive concepts.
@@
Finally, we present our culminating experiment, examining the combined effects of training-time mitigation via SAM and inference-time corrections through UCP. Table 3 demonstrates that integrating SAM-based training with uncertainty-guided interventions nearly recovers the accuracy of cleanly trained models using only five targeted interventions at 20% noise. In stark contrast, models without SAM and using random corrections fall short even after ten interventions. Taken together, these findings underscore the complementary benefit of combining sharpness-aware training with uncertainty-driven interventions, substantially enhancing robustness and preserving interpretability under noisy supervision.

### v2

@@ 1 Introduction
Despite a decade of breakthroughs in deep learning, the inner-workings of end-to-end neural networks remain largely elusive [35, 24, 34]. This lack of transparency stems from their reliance on complex, high-dimensional internal representations, which offer limited insight into their decision-making process [37, 8]. Concept bottleneck models (CBMs) are designed to address this opacity by first predicting a set of human-interpretable concepts, which serve as an intermediate representation for deriving the final decision [19, 3, 38]. For instance, when it comes to the bird species classification problem [54], they first predict the concepts of the bird such as tail shape or body color, and then, infer the bird class based on these attributes, making the decision more understandable (e.g., “this bird is Heermann Gull because it is gray and black in color with an orange beak!”).
Formally, a CBM is trained on triplets of input $x\in\mathbb{R}^{d}$, binary concept vectors $c\in\{0,1\}^{k}$, and output $y$. The concept predictor $g:x\to c$ maps each input to its $k$-dimensional concept vector, and the output predictor $f:c\to y$ consumes those concepts to produce $y$ (see Figure˜1(a)). This two-stage process offers a practical advantage, which is, domain experts can step in and edit the predicted concepts at inference, a process known as concept-level intervention [19]. For example, changing ‘has stripe’ from yes to no, while keeping other concepts such as ‘number of legs’, can shift the prediction from zebra to horse without retraining and can recover much of the lost accuracy. Interventions further support counterfactual analysis [53], allowing users to ask “what if” scenarios by manipulating concepts rather than raw pixels. Such human-in-the-loop interactivity is central to the appeal of CBMs and underscores the need to safeguard concept fidelity.
@@
Specifically, our study reveals that noise significantly compromises key characteristics of CBMs, i.e., interpretability, intervention effectiveness, as well as prediction performance (see Figure˜1 for basic results). We discover that this degradation is primarily driven by a specific set of concepts that are highly susceptible to noise, whose accuracy drop largely accounts for overall performance declines. To mitigate these negative effects, we propose two strategies. First, we utilize sharpness-aware minimization (SAM) [10] during training to build robustness, particularly for noise-sensitive concepts. Second, observing that susceptible concepts often have high predictive uncertainty, quantified by entropy [41], we adopt an uncertainty-based intervention strategy at inference [21, 20, 43]. This strategy prioritizes correcting high-uncertainty concepts, enabling targeted interventions that significantly improve overall task performance. Extensive empirical results validate the effectiveness of this combined approach, consistently recovering the reliability of CBMs despite the presence of noise.
@@ 2 Measuring Impact
In this section, we systematically measure the impact of noise on CBMs across three key dimensions: prediction performance, interpretability, and intervention effectiveness. We begin by outlining our experimental setup (Section˜2.1), followed by an evaluation of the noise effects on prediction performance (Section˜2.2), interpretability (Section˜2.3), and intervention effectiveness (Section˜2.4).
@@ 2.3 Interpretability
Next, we assess the effect of noise on interpretability. As shown in Figure˜2(c), concept prediction accuracy declines from 96.5% in clean conditions to 85.4% with 40% noise on the CUB dataset. Despite appearing modest, the 11.1% drop notably compresses the margin over random guessing in binary classification tasks, where minor concept prediction errors can compound and severely impair target prediction in CBMs. Beyond accuracy, interpretability critically relies on semantic coherence of learned representations. Thus, we measure interpretability using the concept alignment score (CAS) [61], quantifying semantic alignment between learned representations and ground-truth concepts (see Figure˜2(d)). We find CAS decreases substantially with increasing noise, e.g., for each additional 10% increment of noise, CAS drops by approximately 7% on the CUB dataset, indicating severe degradation in interpretability.
Qualitative t-SNE [51] visualizations further illustrate this interpretability decline (see Figure˜4). Clean embeddings for the specific concept ‘blue upperparts’ show distinct, semantically meaningful clusters corresponding to active (i.e., red circles) and inactive (i.e., blue circles) concept instances. However, increasing noise levels result in ambiguous, less separable clusters. Thus, noise compromises the explanatory power and reliability of concepts, diminishing interpretability. Additional examples are provided in Section˜D.12.
@@ 2.4 Intervention Effectiveness
Finally, we investigate how noise limits the utility of post‑hoc interventions, utilizing three concept selection strategies proposed by Shin et al. [43]: (i) Random, which samples concepts uniformly [19]; (ii) UCP (uncertainty of concept prediction), which ranks concepts by predictive uncertainty; (iii) CCTP (contribution of concept on target prediction), which selects concepts exerting the greatest influence on the target prediction. Implementation specifics are detailed in Section˜C.5.
As shown in Figure˜5(a)5(c), the benefits of intervention are increasingly constrained as noise levels rise (i.e., the accuracy curves translate downward along the y-axis). Under the Random strategy, correcting every concept in a model trained with clean supervision yields near-optimal performance at 99.8% accuracy. However, this upper bound progressively declines to 98.4%, 95.7%, 89.4%, and 53.1% as the noise level increases from 10% to 40%. Notably, even exhaustive intervention at 40% noise fails to recover the performance of the clean model without any intervention. Similar trends are observed for the advanced intervention methods (e.g., UCP and CCTP), indicating that the errors induced by noisy supervision in CBMs cannot be rectified through post‑hoc correction alone.
To further evaluate robustness, we deliberately assign incorrect values to selected concepts, emulating the inevitable inaccuracies of expert intervention. As shown in Figure˜5(d)5(e), CBMs trained on clean data tolerate a small number of such mistakes with only marginal performance decline. Under noise, however, even slight intervention errors precipitate pronounced additional loss, revealing a fragility that limits the practical utility of concept‑level correction in realistic human‑expert workflows.
@@ 3 Understanding Mechanisms
In this section, we further analyze how noise affects the internal mechanisms of CBMs, separately examining the concept predictor $g$ and the target predictor $f$. We begin with the observation that noise induces non-uniform degradation of individual concept accuracies (Section˜3.1). We then assess how noise disrupts the alignment between concepts representations and target predictions (Section˜3.2).
@@ 3.1 Non-Uniform Degradation of Concept Accuracy
As shown in Figure˜6, although noise is introduced uniformly across all concepts, accuracy degradation is notably uneven. While the majority of concepts remain comparatively resilient, a distinct minority suffers a disproportionately large decline. We designate this minority as the susceptible concept set, defined as
@@
To better understand this phenomenon, we plot the concept frequency distributions using kernel density estimation (KDE) [5] in Figure˜7. In the clean dataset, concept occurrences are imbalanced, with many low-frequency and a few high-frequency concepts. As noise increases, however, this distribution becomes more uniform; at 40% noise, most concepts occur roughly 2,000 times. This shift may reduce the ability of model to learn rare but informative concepts, as the effective signal-to-noise ratio for such concepts deteriorates. Identifying and understanding these vulnerable concepts is essential, as their degradation can disproportionately affect downstream performance, which we explore in the next section.
@@ 3.2 Disruption in Representation Alignment
We next examine how noise during training reshapes the correspondence between learned representations and their associated target classes. In a noise‑free setting we first identify the five most influential representation dimensions (i.e., concepts) for the class Le Conte Sparrow. These dimensions align with meaningful visual attributes such as ‘white upperparts’ and ‘grey back’. Their importance is quantified by the absolute magnitude of the weights in the linear target predictor $f$, a widely used proxy for feature saliency [25, 22, 6]. We then trace how these saliency values change as the proportion of corrupted labels increases.
As illustrated in Figure˜8(b), when the noise level is moderate (i.e., 20%) the relative ordering of the predictive dimensions is largely preserved. At a higher noise level (i.e., 40%), however, their saliency declines markedly. Informative dimensions lose influence, whereas irrelevant or spurious ones gain disproportionately large weights (see Figure˜8(c)). Specifically, dimensions linked to ‘grey back’ or ‘yellow upper tail’ receive minimal weight, while a related attribute such as ‘orange upper tail’ is assigned a negative weight. These shifts indicate a pronounced misalignment between the learned representations and their intended semantics.
We further relate this effect to concept-level susceptibility. In Figure˜6, its five most influential dimensions, indicated with red arrows, coincide with the concepts most susceptible to noise. The same pattern emerges across the full dataset. Among 200 classes, 189 display an exact overlap between the five most influential dimensions and the noise-susceptible set. This widespread overlap exposes a critical weakness, i.e., when concepts that are noise-sensitive coincide with distorted representations, performance degrades sharply. Thus, the intersection of susceptibility and disrupted alignment emerges as a principal source of instability, emphasizing the need for robust training methods to prevent representation collapse under noisy supervision.
@@ 4 Mitigating Effects
We have shown that noise severely degrades the key aspects of CBMs, primarily due to a subset of concepts (i.e., susceptible concept set) that are especially vulnerable to noisy supervision and frequently coincide with representational misalignments. To address this, we introduce two mitigation strategies that operate at different stages of the model lifecycle: during training and inference. First, we incorporate sharpness-aware minimization (SAM[10] at training time to improve the prediction accuracy of susceptible concepts (Section˜4.1). Then, at inference time, we apply an uncertainty-based intervention strategy that prioritizes corrections on unreliable concept predictions (Section˜4.2).
@@ 4.1 Training-Time Robustness via SAM
where the inner maximization is approximated by $\widetilde{\bm{\varepsilon}}(\bm{w})=\rho\frac{\nabla_{\bm{w}}\ell_{\mathrm{train}}(\bm{w})}{\|\nabla_{\bm{w}}\ell_{\mathrm{train}}(\bm{w})\|_{2}}$. Gradient updates are then computed at $\bm{w}+\widetilde{\bm{\varepsilon}}(\bm{w})$. By steering solutions toward flatter optima, SAM attenuates the performance degradation caused by noisy labels and other small perturbations [10, 2], rivaling the best dedicated noise-robust training methods [17, 62, 1].
@@
[…]
@@
We compare SAM-trained CBMs with SGD-trained ones (i.e., Base) across concept accuracy, task accuracy, and CAS on the CUB and AwA2 datasets under varying noise levels. As shown in Table˜1, SAM consistently improves all metrics across noise conditions. Notably, even modest gains in concept accuracy lead to substantial improvements in downstream task performance. For instance, under 40% noise in AwA2, a 0.60% increase in concept accuracy leads to a 4.68% gain in task accuracy. Previous work [2] attributed the robustness of SAM against noise to its $\ell_{2}$-regularization on final layer weights and intermediate activations. In Appendix˜A, we extend the theoretical analysis of Baek et al. [2] to CBMs, demonstrating that the same regularization effect holds.
We also emphasize that SAM is particularly effective at recovering concepts in the susceptible concept set. Table˜2 shows that while concept accuracy in the non-susceptible set remains stable, SAM yields substantial gains in the susceptible set across all noise levels, e.g., yielding gains of +3.85% and +4.07% under 20% and 40% noise, respectively. These findings underscore the effectiveness of SAM in selectively enhancing the learning of noise-sensitive concepts, thereby enhancing both predictive robustness and model interpretability under corrupted supervision.
@@
Figure 9: Intervention results comparing baseline and susceptibility-based corrections. The dotted line indicates the clean-model accuracy without intervention.
@@ 4.2 Inference-Time Robustness via Uncertainty-Guided Intervention
Although training with SAM effectively mitigates the adverse effects of noise, additional performance gains are achievable through targeted inference-time interventions. Motivated by our earlier findings that only a small subset of concepts significantly contributes to performance degradation, we prioritize corrections based on susceptibility, defined as the magnitude of accuracy decline under noise. Specifically, we select the five most susceptible concepts per class for correction. As shown in Figure˜9, susceptibility-guided corrections significantly outperform baseline interventions across varying noise levels. Remarkably, correcting even a single highly susceptible concept can yield accuracy improvements of around 10% at higher noise intensities (e.g., 20% and 30%). Moreover, fewer than five targeted interventions frequently achieve or surpass the accuracy of models trained on clean data. These findings underscore that noise disproportionately affects a small subset of highly predictive and noise-sensitive concepts, and that targeted correction of these concepts substantially restores overall model performance.
@@
While susceptibility-based concept selection is a highly effective intervention strategy, it is not feasible to perform in realistic scenarios since it requires the access to a model trained with clean labels. To circumvent this challenge, we explore predictive uncertainty as a practical surrogate for susceptibility. Intuitively, concepts that contribute little to target prediction or are noise-sensitive should exhibit elevated uncertainty. To validate this hypothesis, we measure the Pearson correlation between concept-level uncertainty (quantified by the entropy of their predictive distributions [41]) and susceptibility. As depicted in Figure˜10, we observe a positive correlation within the subset of susceptible concepts, whereas no consistent relationship emerges among non-susceptible concepts. These results suggest that predictive uncertainty effectively, though not definitively, reflects susceptibility induced by noise, supporting its use as a practical criterion for inference-time interventions. In Appendix˜B, we provide theoretical justification demonstrating that, under reasonable assumptions, uncertainty-based selection asymptotically approximates susceptibility-based concept selection.
@@
[…]
@@
[…]
@@
Finally, we present our culminating experiment, examining the combined effects of training-time mitigation via SAM and inference-time corrections through UCP. Table˜3 demonstrates that integrating SAM-based training with uncertainty-guided interventions nearly recovers the accuracy of cleanly trained models using only five targeted interventions at 20% noise. In stark contrast, models without SAM and using random corrections fall short even after ten interventions. Taken together, these findings underscore the complementary benefit of combining sharpness-aware training with uncertainty-driven interventions, substantially enhancing robustness and preserving interpretability under noisy supervision.
5 Discussion
On the independence assumption
Our primary objective is to provide a systematic analysis of how noise affects the performance and interpretability of CBMs. To isolate basic effects and to evaluate mitigation strategies in a controlled setting, we sought an experimental design in which concept noise and target noise can be manipulated independently. This design allows us to vary each factor without confounding and to attribute observed changes to a single source of corruption. For this reason, we considered the independence assumption a reasonable and transparent starting point for an initial analysis across a range of noise levels.
Nonetheless, we acknowledge that in real-world data, concept and target noise may be correlated. To account for this, we include additional experiments under more realistic conditions, simulating a correlated noise scenario where concept corruption occurs only when the target label is flipped, thereby inducing dependence between concept and target noise. The corresponding results are provided in Section˜D.4. In summary, even under dependent noise, the positive association between uncertainty and susceptibility persists for susceptible concepts. This pattern remains consistent across noise levels and indicates that our analysis and mitigation approach are effective in more realistic correlated noise settings. However, as the dependent noise considered here is still relatively simple, it is important to account for more complex correlation structures that may arise in real-world data.
Benign appearance of target noise
In our experiments, we observed that CBMs are highly sensitive to concept noise but relatively robust to target noise. This robustness can be attributed to the structure of the CBM, where the target predictor $f$, following prior work [19], is implemented as a single linear layer mapping concepts to the target. Its limited capacity makes it less able to overfit noisy target labels, reducing sensitivity to target corruption. Moreover, CBM training primarily burdens the concept predictor $g$. When $g$ produces stable, accurate concept representations, $f$ operates on relatively clean inputs, so even severe target noise has a muted effect on final accuracy.
Generalization across CBM variants
We primarily analyzed the impact of noise on CBMs using the most representative formulation, the original CBM [19]. To assess whether our methodology, particularly the application of SAM, generalizes to other CBM variants, we extend our analysis accordingly. Specifically, we evaluate several representative variants (e.g., CEM [61], ECBM [59], AR-CBM [14], and SCBM [52]) to investigate how the presence of noise influences their performance. As summarized in Section˜D.2, all variants exhibit clear degradation in both concept and task accuracy as the noise level increases.
We further investigate whether the ability of SAM to mitigate the susceptible sets generalizes across different CBM variants. For SCBM, SAM showed clear improvements over SGD: at 20% noise, 75.8 vs. 74.9 (+0.9), and at 40% noise, 60.1 vs. 56.4 (+3.7). For AR-CBM and CEM, the gains were smaller: AR-CBM at 20% noise 71.9 vs. 71.3; at 40%, both 53.9; CEM at 20%, 73.3 vs. 73.1; at 40%, 54.6 vs. 54.5. These results indicate that the benefits of SAM vary across CBM variants and that resilience to susceptible concepts may depend on the underlying model design. While our focus remains on CBMs, we acknowledge that developing a broader and systematic framework encompassing multiple variants is important for understanding generalization under noisy conditions.
Generalization across modalities
CBMs are now being explored beyond vision, including settings that combine CBMs with large language models [49, 46]. The core CBM process first predicts concepts and then predicts the final target, which does not depend on a specific modality. Recent LLM based CBM studies employ a similar two stage design, so we expect our main findings to transfer to these settings. That said, concept annotated datasets for LLM are not yet systematically established. An important next step is to curate well-defined concept annotated datasets suitable for LLM applications and to analyze how label noise in such datasets affects CBMs.
Rationale for employing SAM
We adopted SAM based on prior findings by Baek et al. [2], which demonstrate that SAM is a representative and often effective approach for alleviating label noise, and it can outperform alternative methods [62, 1, 17]. In our experiments, SAM also mitigated the impact of noisy concept annotations in CBMs, particularly for susceptible concepts, confirming its suitability as a mitigation method in this setting. We do not claim that SAM is the only solution; rather, we present it as one effective approach to the observed problem and note that identifying other suitable mitigation methods for different CBM variants is also important.